### v2

# Bridging Developer Needs and Feasible Features for AI Assistants in IDEs

Abstract
Despite the increasing presence of AI assistants in Integrated Development Environments, it remains unclear what developers actually need from these tools and which features are likely to be implemented in practice. To investigate this gap, we conducted a two-phase study. First, we interviewed 35 professional developers from three user groups (Adopters, Churners, and Non-Users) to uncover unmet needs and expectations. Our analysis revealed five key areas: Technology Improvement, Interaction, and Alignment, as well as Simplifying Skill Building, and Programming Tasks. We then examined the feasibility of addressing selected needs through an internal prediction market involving 102 practitioners. The results demonstrate a strong alignment between the developers’ needs and the practitioners’ judgment for features focused on implementation and context awareness. However, features related to proactivity and maintenance remain both underestimated and technically unaddressed. Our findings reveal gaps in current AI support and provide practical directions for developing more effective and sustainable in-IDE AI systems.
Index Terms:
Human-AI Collaboration, Artificial Intelligence, Generative AI, LLMs, Intelligent Assistants, Integrated Development Environment, Programming
I INTRODUCTION
The increasing integration of Artificial Intelligence (AI) tools in software development is transforming the way developers interact with their Integrated Development Environments (IDEs) [1, 2, 3]. These AI-powered assistants facilitate tasks like code generation, debugging, and optimization with the goal of boosting productivity and reducing the cognitive load associated with repetitive programming tasks [4, 5, 6, 7]. As AI-based systems evolve rapidly, teams that build tools for tech creators face two key challenges: understanding how coding specialists actually use AI inside their IDEs and deciding which features are worth investing in.
However, actionable insights into in-IDE human-AI experience (HAX) are still limited. While recent studies have explored developers’ general expectations for AI-assisted tools, few focus specifically on in-IDE interactions, particularly the successes and challenges developers face and their needs with these tools [7]. Even fewer studies consider differences across user groups: those who adopted AI into their workflow, those who abandoned it, and those who have never tried it. These distinctions are important for informing development and marketing strategies [8]. Moreover, these studies often focus on usage patterns or user perceptions without connecting them to concrete feature development decisions. As a result, we know little about how specific user needs translate into buildable and valuable AI features.
To address existing knowledge gaps, we conducted a multi-method study in two phases: qualitative semi-structured interview [9] and prediction market [10]. First, we interviewed 35 tech creators across three usage groups (Adopters, Churners, and Non-users) to gain qualitative insights into their experiences, challenges, and unmet needs with current AI features. This phase reveals unmet needs and expectations across five areas: Technology Improvement, Interaction, Alignment, Simplifying Skill Building, and Programming Tasks. Yet, understanding user needs alone does not directly translate into feasible product development strategies. Prioritizing feature development requires additional insight into both technical feasibility and organizational readiness. To achieve this, we complemented our qualitative findings with a prediction market, an approach in which participants make predictions on future events, leveraging collective judgment to estimate outcomes. We asked 102 internal practitioners to predict the likelihood of implementing 12 user-desired features. This allows us to integrate user-driven insights with practical considerations from the very people who shape our product development. Our results reveal that internal practitioners correctly anticipate the implementation of features around AI in Implementation and Context-aware Technology. In contrast, features requiring proactive behavior and maintenance support remain undervalued and underdelivered. These findings highlight a misalignment between developer needs and current feasibility assessments, pointing to concrete opportunities for more helpful and forward-looking AI assistance in IDEs.
With this work, we aim to provide industry practitioners and researchers with structured evidence-based insights to guide product decisions. Specifically, offer:
  • A categorization of user needs for in-IDE HAX to outline the technological improvements needed to make AI tools more useful;
  • A comparison of perspectives across Adopters, Churners, and Non-Users, revealing group-specific concerns that influence adoption and retention, such as trust and technical issues among Churners and ethical reservations and skill development concerns among Non-Users.
  • A mapping of the externally voiced needs to internal perceptions of feasibility, identifying areas of alignment (e.g., implementation and context-awareness) and areas of mismatch (e.g., proactivity and maintenance support).
We are the first to triangulate detailed user needs with internal feasibility estimates using prediction markets. This approach enables organizations to better prioritize AI features grounded in what users want and what teams can build.
II BACKGROUND
AI-powered programming assistants have become increasingly common, with tools such as GitHub Copilot, Cursor, and JetBrains AI integrated into developer workflows [4, 5, 6]. While their technical capabilities continue to evolve, research into how developers experience and evaluate such tools has also grown.
Recent studies have investigated what professional developers need and expect from AI-powered programming assistants [11, 12, 13, 14, 15, 16, 17]. Across contexts, developers consistently seek tools that reduce repetitive effort and provide relevant suggestions without introducing new debugging burdens [11, 16, 15]. Effective assistance is expected to align with the current task, understand code context, and support natural language interaction [11, 13, 18].
Transparency and control are also emphasized, including the ability to steer and correct outputs and control context [12, 11]. Several studies highlight that experienced users tend to expect more than raw generation [13, 12]. For example, tools are valued when they adapt to user intent and surface relevant, task-specific information [11, 13, 19].
At the same time, integration challenges persist. Developers often report friction when tools disrupt their workflow or fail to fit specific environments [14, 15]. Outside of initial implementation stages, non-functional requirements such as performance and scalability may be harder to maintain [14]. These insights point to the need for assistants that can flexibly support diverse roles, offer customization, and operate within existing development processes [12, 14, 11].
Despite these insights, most studies examine developer interaction with AI in broad or external settings, without addressing the unique characteristics of IDE-based workflows. This distinction is important. According to large-scale industry reports, up to 97% of developers rely on IDEs for software tasks [20]. In-IDE AI assistants operate within a structured environment that offers richer contextual signals such as file scope, project layout, and cursor history. These tools have the potential to deliver more context-aware support while minimizing context switching. Yet, their design introduces new constraints, including the need to remain non-intrusive and responsive in real time. Therefore, in-IDE AI tools shape distinct user needs, expectations, and challenges that warrant focused investigation.
Moreover, existing studies tend to treat developers as a single user group, overlooking differences across adoption stages [7]. However, industry practice suggests that design and feature prioritization must account for the heterogeneous needs of distinct groups: Adopters, Churners, and Non-Users [8, 21]. These groups not only engage (or not) with AI tools differently, but also hold divergent mental models, trust levels, and expectations, which may lead to different product requirements or marketing strategies. Recognizing these “personas” is necessary to design AI assistants that support retention, re-engagement, or onboarding.
At the same time, even when user needs are clearly articulated, translating them into product decisions remains a challenge. Feature roadmaps must also reflect internal constraints such as technical feasibility, resource allocation, and infrastructure readiness. Traditional prioritization techniques, such as expert panels and surveys, often fall short by providing limited collective insights or biased individual opinions [22]. An alternative approach, prediction markets, has been shown to effectively aggregate collective expert judgment to forecast the likelihood of specific outcomes [10]. A prediction market is a structured method wherein participants buy and sell predictions about future events, effectively leveraging the “wisdom of the crowd” to estimate probabilities accurately [10, 23].
Thus, this paper combines qualitative interviews and prediction markets, integrating rich user insights with realistic assessments from internal stakeholders. This dual perspective enables a grounded understanding of both user needs and internal feasibility, offering a practical contribution to the design and deployment of in-IDE AI assistants.
III UNDERSTANDING DEVELOPER NEEDS
In this section, we examine how developers experience AI inside their IDEs when coding. Our goal is to provide a structured understanding of what developers need from AI assistance, what challenges they face, and how these vary across adoption groups.
We address the following research questions:
  • RQ1. What are developers’ concrete needs when using AI tools within IDEs?
  • RQ2. How do these needs differ between Adopters, Churners, and Non-Users?
III-A METHOD
The study was conducted in line with our institutional ethical standards, adhering to the values and guidelines outlined in the ICC/ESOMAR International Code [24].
III-A1 Recruitment
We recruited participants from users of JetBrains, a software engineering company that also develops AI-based tools for software development, to ensure a consistent experience with development environments. Note that with that, participants could use any in-IDE AI assistant. Invitations were sent via email to consenting subscribers and promoted through the authors’ social media.
Respondents completed a qualification survey designed to assign them to one of three groups: Adopters, Churners, and Non-Users. We chose these three groups to capture various experiences with AI tools in IDEs [8]. These groups are referred to using the prefixes A, C, and N, respectively, in participant identifiers throughout the paper.
In the survey, we collected data on job roles, experience levels, IDEs used, AI tools tried for coding, and duration of AI usage. Out of 380 responses, we iteratively invited participants until data saturation was reached [25], which means that no new themes or insights emerged from the last interviews. As a result, we conducted 35 interviewees: 15 Adopters, 12 Churners, and 8 Non-Users.
Most of the participants were developers (32) with overlapping roles: 9 team leads, 8 architects, and 5 DevOps engineers. The sample included 28 seniors, 5 mid-level practitioners, and 2 juniors, with experience ranging from under 5 years (9) to over 11 years (16). Participants chose a reward between a USD 100 Amazon gift card or an equivalent-value pack of all JetBrains products.
III-A2 Data Collection
The interviews followed a semi-structured script developed by the first two authors, reviewed by certified UX researchers at JetBrains, and piloted with four internal participants [9]. The script included open-ended questions on AI experience, utilized tools, challenges and successes with in-IDE HAX, unmet needs, desired features, and visions for AI in coding.
To account for differing usage patterns, we tailored the script per group. For Adopters, the questions focused on their ongoing use of AI tools, exploring typical tasks, workflows, challenges, and successes. Churners reflected on their initial adoption and further disengagement, identifying reasons for discontinuation and missed opportunities. Non-Users were asked about their reasons for not adopting AI, any potential barriers and possible motivations that could encourage them to begin using these tools. This ensured group-specific depth while maintaining thematic consistency across interviews.
III-A3 Data Analysis
All interviews were transcribed using Condens [^1] Condens: a user research repository for analyzing qualitative data. https://condens.io/. We applied a thematic analysis [26, 27] to identify user-reported successes, challenges, unmet needs, and future expectations for in-IDE AI.
The first two authors reviewed the first transcript together to agree on the analysis and the code-deriving process. They then went into an iterative inductive process, where they independently assigned Codes to quotes from the transcripts. Eventually, they compiled a list of 161 codes, met again to add or remove codes, and resolved any disagreements until full consensus was reached. The final versions of the codes were then assigned to the corresponding quotes. The codes were then grouped into Thematic groups representing key areas of need. These were organized into broader Topics that describe the categories of in-IDE HAX needs.
For example, the following quote:
“The most useful thing you could do with AI would be a summary of a Pull Request.” — N13
was coded as “Code Review Support”, assigned to the group “Code Optimization”, under the topic “Simplifying Programming Tasks”.
All materials used for data gathering, together with a full demography table of participants and a summarized codebook, are included in Supplementary Materials [28]. Full transcripts of interviews cannot be shared due to the company’s confidentiality policy.
III-B FINDINGS
Figure 1: The Categorization of Users’ Needs in HAX Within IDEs With Descriptions
In Figure 1, we present our categorization of developer needs regarding Human-AI Experience within IDEs, based on a thematic analysis of 35 interviews. It includes five overarching topics: Technology Improvements, Simplifying Programming Tasks, Technology Interaction, Technology Customization, and Simplifying Skill Building.
Technology Improvements
Participants stressed the need for more reliable AI systems that reflect the latest advancements of the field. Participants also raised concerns about the exhaustion of training data, the perpetuation of outdated or erroneous information, and the plateauing progress of AI, all of which could undermine the effectiveness and accuracy of tooling. Participants also noted current shortcomings in output quality and emphasized the importance of thoroughly reviewing and validating AI-generated code to avoid overreliance.
“If I felt the AI system was really reliable, it would have saved me time because I wouldn’t need to double-check and could just apply the right solution right away.”C159
Privacy and security were also identified as critical requirements for AI adoption. Participants favored on-premise or local AI solutions to ensure data confidentiality and comply with organization-specific or national restrictions. They stressed the importance of preventing vulnerabilities and maintaining the integrity of AI-generated code, including making such code clearly identifiable to facilitate inspection and verification.
“Regulations restrict us from using AI when it involves sending data to the cloud. My company is particularly interested in on-premise language models to maintain control and keep data within the country or even company.”A77
Context awareness was seen as essential for generating relevant and accurate code. The participants highlighted the need for mechanisms to manually set, view, and edit the context provided to the AI assistant. Some expressed frustration over the lack of control in current systems and noted that greater transparency in how context is constructed could also help mitigate privacy concerns.
The participants also called for better alignment of AI tools with their existing development practices. They praised models that generate suggestions consistent with the project structure and coding conventions, such as recommending appropriate libraries or learning from examples. At the same time, they noted issues with overly verbose, redundant, or mismatched output that diminished the value of AI-generated assistance.
Proactive features were seen as an opportunity to shift AI from a reactive assistant to a more autonomous collaborator. Participants envisioned systems that monitor ongoing work, detect issues, and make targeted improvements, such as flagging performance bottlenecks or managing dependencies, as long as these actions remain in the trusted IDE environment and do not involve external model training.
“It would be cool if it could search for problems along the way. It could say ‘Did you realize you just created an out-of-bounds condition?’ ”A44
Accessibility was another recurring concern. The interviewees pointed to the need for tools that are affordable and adaptable across infrastructural, legal, and experiential boundaries. Suggestions included natural language programming, universal language support, and modular pricing models that lower the barrier to entry, particularly in regions with limited connectivity or restrictive regulations.
The participants also emphasized the importance of ethical safeguards. They called for clear rules and restrictions to guide responsible AI use and suggested that automated mechanisms may be necessary to ensure compliance and accountability,particularly as AI becomes more integrated into daily development workflows.
“AI, like anything with the potential to destroy jobs, carries a huge responsibility. I really hope that the community and organizations invest heavily in getting it right. Otherwise, recovering might be really difficult.”A37
Simplifying Programming Tasks
The participants emphasized the potential of AI to reduce manual effort by supporting routine and repetitive activities across the software development lifecycle. By automating low-level tasks during ideation, implementation, testing, maintenance, and optimization, developers could shift focus to more complex and creative work. Many participants envisioned AI evolving into an active collaborator capable of supporting ideation, implementation, testing, maintenance, and optimization.
“I would say it should be like a colleague, someone who understands what I wrote and can guide or help me where I lack knowledge.”A26
A common expectation was for AI to handle well-defined repetitive work, such as template-based coding, documentation generation, and routine debugging. This division of labor was seen as essential for maintaining developer productivity without increasing cognitive load.
The interviewees also emphasized the importance of AI in supporting codebase navigation. In large or unfamiliar projects, features such as database analytics, diagram generation, and in-context code explanations were considered particularly valuable. These capabilities were viewed as instrumental in reducing ramp-up time and enabling a more seamless understanding of system architecture.
“If you gave the AI the company documentation and code and then let it automatically detect what’s needed, this would help newcomers follow the company’s standards.”C94
The participants also expressed interest in AI tools that integrate IDE with external systems, such as CI/CD pipelines, version control platforms, and support ticketing systems. The ability to access and coordinate AI assistance across these environments was seen as a way to increase traceability and streamline workflows spanning multiple tools.
“Right now, so many things come through email, then there’s the support ticketing system, and, of course, our Git repositories. There could be ways to integrate all these even better.”C118
Technology Interaction
The interviewees emphasized the importance of smooth internal system integration, where AI tools interact efficiently with other features within the IDE. Integrating AI into the IDE, especially through embedded chat functionality, was considered particularly beneficial, reducing context switching and, therefore, avoiding unnecessary interruptions.
“I don’t need to switch to a different tab or window. I don’t have to go through that whole process, so that’s definitely very helpful. It saves time and makes the process more efficient.”C122
The participants expressed a desire for other types of uninterrupted experiences that allow them to interact with AI in flexible ways, including via voice commands or wearables such as Neuralink. They also showed interest in features like inter-AI interaction, in-line code explanations, and customizable key mappings to further streamline workflows.
At the same time, several challenges were reported. These included slow response times, frequent logouts, and limitations in chat functionality, such as copy/paste issues or dead-ends in conversations where the AI failed to respond. Participants also noted struggles with reproducibility in code completion, UI inconsistencies, and interference between suggestions from different completion providers. Such issues were seen as disruptive to overall workflow continuity.
  • *Note: Symbols represent the participant groups: Adopters (⚫), Churners (★), and Non-Users (▲). Presence of a symbol indicates that group members mentioned the corresponding theme in the associated interview section.
Figure 2: Comparative Analysis of Thematic Group Mentions Across User Types.
Technology Customization
The interviewees stressed the importance of adjustable AI systems that offer a high degree of personalization and flexibility, allowing users to retain control as decision-makers. Key features included configurable AI, the ability to choose between different models, and the option to customize AI actions based on specific tasks or workflows. Participants also expressed a desire for task-specific setups that allow AI to adapt to unique user preferences and working styles.
To support workflow optimization, participants were also interested in tools that could help manage distractions, track time, and summarize daily activities within their development ecosystem. The ability to toggle time tracking, manage task tickets, and maintain a personal prompt library was highlighted as essential for staying organized and improving productivity.
“At my job, I have around 20 tickets pending, and sometimes I forget which one I’m currently working on and what to do next, so I need something that can help me stay organized.”N14
Simplifying Skill Building
The participants noted that AI tools often have a steep learning curve. They identified the need for better user education, particularly around effective prompt formulation. A shareable prompt library was seen as a possible valuable resource, as many reported difficulties in crafting prompts that yield useful results. Additionally, the unfamiliar interface of some AI systems contributed to these challenges. Several Non-Users mentioned that the primary reason they had not adopted AI tools was the time required to become proficient.
“It’s a useful tool if employed correctly. But there is a learning curve. If I figure out how to work with it efficiently, how to set up the right prompts, etc., it should speed up my work.”N84
The participants also emphasized the role of AI as a learning enabler and mentor, helping developers acquire new skills and understand new concepts. Features such as step-by-step output reveal were considered helpful in fostering long-term skill development through more interactive and adaptive guidance.
“This way, you wouldn’t have to keep requesting new hints would be more interactive, allowing you to reveal the steps as you go and adjust based on what you need.”A58
Answer to RQ1. Developers emphasized the need for accessible, proactive, reliable, secure, and context-aware technology that adapts to their workflows and remains ethical and up-to-date. They seek AI tools that support various aspects of development, from code generation and testing to codebase navigation, focusing on non-interruptive, customizable, and learning-supportive interfaces. Moreover, developers prioritize smooth internal integration of AI with existing systems, emphasizing the need for intuitive user education and skill development to maximize AI’s potential within their workflows.
Comparative Analysis of Codes Among Groups
We present a comparative analysis of the Thematic groups among Adopters, Churners, and Non-Users in Figure 2, focusing on participants’ success stories (or anticipated advantages for Non-Users), challenges (or barriers for Non-Users), and needs in the context of in-IDE HAX. Lastly, we explore the similarities and differences at the Codes’ level.
  • Similarities Across All Groups:
    Concerns about AI displacing jobs were widespread, reflecting apprehension about job security and the need for AI tools to complement rather than replace human capabilities. Participants expressed a general interest in AI’s role in code generation, explanation, and evaluation, recognizing its potential to simplify coding tasks. We identified the lack of sufficient project-wide context awareness in AI tools as a common challenge, with participants calling for improvements in this area to boost AI’s effectiveness.
    Ethical and privacy concerns, such as plagiarism and data protection, were also frequently mentioned, emphasizing the need to address these issues to build trust. The participants also noted the risk of overreliance on AI, highlighting the importance of balancing AI assistance with personal skill development. Despite these concerns, there was a shared sense of optimism, with many participants acknowledging the ongoing improvement of AI tools over time.
  • Similarities Between Adopters and Churners:
    Both groups faced challenges with AI integration, such as code completion clashes and frequent logouts, confirming that such technical difficulties negatively affect user retention and satisfaction. There was a shared interest in AI functionalities that proactively assist with debugging, optimization, and documentation, stressing a demand for more advanced AI capabilities. While some Adopters and Churners reported similar challenges when using AI assistants, Adopters were more likely to persist through issues, experiment with workarounds, and continue using the tools despite their limitations. This contrast suggests that factors beyond the AI’s technical performance influence user retention. Further research is needed to understand what shapes developers’ willingness to tolerate imperfections and continue engaging with AI tools over time.
  • Differences Between Adopters and Churners:
    Adopters demonstrated a more positive outlook, frequently expressing optimism and enthusiasm. They also reported interest in advanced features like Inter-AI Interaction and Cross-Model Prompting, reflecting a deeper integration of AI into their workflows. Additionally, Adopters showed a clear preference for customization and personalization of AI tools to fit their specific needs.
    In contrast, Churners often exhibited skepticism and dissatisfaction. They emphasized security and privacy concerns more, favoring on-premise AI solutions due to trust issues with cloud-based services. Churners also highlighted technical limitations such as slow response times, outdated data, and limited creativity from language models, which contributed to their discontinuation of AI usage. It is worth noting that several Churners were forced to stop their AI usage due to company or country restrictions. They mentioned being overall satisfied with their experience and emphasized that more secure and on-premise solutions could get their companies to reconsider the restrictions.
  • Unique Themes Among Non-Users: Despite mostly being aware of AI capabilities, Non-Users had not adopted these tools, indicating potential barriers to initial adoption. A strong focus on personal skill development emerged among Non-Users, suggesting a preference for traditional learning methods over AI assistance. Moreover, Non-Users expressed unique ethical concerns about the environmental impact and ownership issues, highlighting the influence of broader societal implications on their decision not to engage with AI technology.
Answer to RQ2. Adopters show enthusiasm for advanced AI features and customization, suggesting that further innovation can strengthen their engagement. Churners, however, highlight technical issues and trust concerns. This indicates the need for enhanced reliability and privacy-preserving measures. Non-Users emphasize personal skill development and ethical concerns and highlight a barrier to adoption that can be mitigated with better education and clearer ethical guidelines.
IV ESTIMATING FEATURE FEASIBILITY VIA PREDICTION MARKET
Table I: Prediction accuracy and market activity across tools for proposed features.
Feature Tool Resolved Acc.? Prob. Volume Liq.
Multi-file Generation Any Tool 98% 1946 570
JetBrains IDE 92% 4288 610
Project-wide Context Awareness Any Tool 82% 440 290
JetBrains IDE 41% 2238 790
Sketch2Code Any Tool 90% 1019 310
JetBrains IDE 10% 500 270
Proactive Documentation Any Tool 33% 878 310
JetBrains IDE 17% 1115 370
Proactive Automation Any Tool 24% 590 290
JetBrains IDE 13% 440 270
Annotating AI-Generated Code Any Tool 18% 733 290
JetBrains IDE 18% 640 270
Business-context Awareness Any Tool 22% 272 150
JetBrains IDE 10% 190 130
Daily Ecosystem Summarizer Any Tool 15% 295 190
JetBrains IDE 14% 510 250
AI Real-Time Refactoring Any Tool 19% 950 370
JetBrains IDE 8% 420 250
Proactive Optimization Any Tool 19% 650 310
JetBrains IDE 8% 325 210
Proactive Debugging Any Tool 6% 750 210
JetBrains IDE 8% 340 150
Brain-computer Interface Coding Any Tool 4% 1241 490
JetBrains IDE 1% 2150 510
  • *Note: Each row shows a market prediction about a feature launch. Resolved shows if the feature was launched (✓) or not (✗). Acc.? indicates if the prediction was correct. Prob. is the final market probability. Volume and Liq. reflect participant engagement.
In the first phase of the study, described in Section III, we explored the needs of different groups of users in regards to AI. In the subsequent phase, we investigated how the market responds to these needs in a short-term period. We were also interested in whether the tech creators’ community can successfully predict which of the needs are likely to be met. Therefore, we conducted a prediction market “tournament” on JetBrains’ internal platform.
Prediction markets are exchange-traded markets where people’s bets, with payoffs depending on a future event, can be used as a measure of the common beliefs. The prevailing price of the bet represents the collective belief in the probability of that event. This effect follows from the principle that aggregating many independent bets yields accurate forecasts, the so-called wisdom of the crowd [10]. Previously, prediction markets have been successfully used to forecast a broad range of outcomes from presidential votes [29], to replicability of scientific research [30], and the future of real products on the market [31]. In our research, through prediction market bets, we examined which user needs and feature directions tool builders recognize and consider relevant. We were also interested in testing whether these bets reflect an accurate intuition about how the field is likely to evolve, measured by whether the predicted features are eventually implemented or not. Consequently, our research question for Phase 2 is
  • RQ3. To what extent do practitioners recognize and correctly anticipate which AI features inspired by user needs are likely to be implemented, as reflected in their prediction market activity?
All materials used for data gathering at this phase, as well as the full demography table and a summarized report of the prediction market, are in Supplementary materials [28].
IV-A METHOD
IV-A1 Materials
Based on insights from interviews, we generated an initial list of 38 potential AI-based IDE features [28]. This list was reviewed by seven experienced researchers from industry and academia: three specialists in human–AI experience, one assistant professor in computer science, and three senior software engineering researchers. All reviewers had more than five years of professional experience. Through this collaborative process, we selected 12 features for the prediction market. These were chosen based on their perceived market feasibility within the next year and the possibility of defining each feature with a high degree of clarity.
Based on the final list of features, we formulated 24 binary questions for the prediction market. Each question asked whether a given AI-based feature would be released within the next six months, either in JetBrains IDEs or any other AI-enabled development tool. These questions were published on JetBrains ’ internal prediction market platform, where employees place virtual bets to estimate the likelihood of future events. All 24 questions were grouped into a single tournament, where cash prizes (\$1500, \$1000, \$500) were to be awarded to the top three predictors on the leader board to incentivize participation.
IV-A2 Participants
All participants of the prediction market were employed by JetBrains. A total of 102 participants [28] across 84 teams placed bets on at least one question during the tournament. The median tenure was 4 years and 10 months. The most common roles included Software Developer (34 participants), Team Lead (20), Product Manager (6), Analyst (6), QA Engineer (5), and ML Engineer (4).
IV-A3 Data collection
The tournament was launched in December 2024. Each participant received 500 virtual coins to place bets on the outcomes of 24 binary questions. For each question, participants could allocate any number of coins to the YES or NO outcome. Before trading, all participants were informed about tournament rules, trading procedures, and strategies for making informed bets. During trading, the current market probability for each question was visible. Bidding closed on June 1, 2025, and all questions were resolved in July 2025. Participants who made correct predictions earned coins, and the top performers were awarded monetary prizes.
The final resolutions were made by a panel of four expert judges, each with over seven years of professional experience in software engineering, AI/ML research, and product development at the company. The panel of judges did not participate in the initial preparation of the features for prediction. Before resolution, the organizing team compiled a summary report [28] on the market status of each feature based on analysis of the feature set of JetBrains’ IDEs, and 20 other tools that were considered, e.g., Windsurf, Cursor, GitHub Copilot. This document served as a reference for the judges, who reached each final decision through several rounds of structured discussion.
IV-A4 Data analysis
We analyzed the prediction market results along three dimensions: participant engagement, accuracy of forecasts, and final probability estimates. Engagement was measured using two metrics: trading volume and market liquidity. Trading Volume captures the total number of virtual coins bet across both outcomes of a question, reflecting overall activity. Market Liquidity reflects the breadth of participation and was calculated by assigning a base value of 50 to each question, with 20 additional points added for every unique participant who placed a bet. Accuracy was assessed by comparing the community’s majority prediction with the actual outcome, as defined by the resolution criteria. The final probability estimate for each question was calculated as the proportion of coins placed on the YES outcome relative to the total of coins bet on that question.
IV-B FINDINGS
To assess whether internal practitioners recognize and correctly anticipate which AI features inspired by user needs are likely to be implemented, we analyzed prediction market outcomes along three dimensions: forecast accuracy, final probability estimates, and participant engagement. The results are summarized in Table I.
Forecasting performance
Ultimately, experts concluded that 6 of the 12 features have appeared on the market in any coding tool, and 2 of the 12 appeared specifically in JetBrains IDEs. The market produced accurate forecasts for 20 of the 24 questions (83%). Accuracy was higher for internal JetBrains releases (92%) than for external tooling (75%), suggesting better visibility into internal development plans and greater uncertainty around the broader field.
All four incorrect forecasts were false negatives: participants assigned low probabilities to features that were actually released. These included Project-wide Context Awareness in JetBrains IDEs, and three features that appeared in other tools: Annotating AI-Generated Code, Business-context Awareness, and the Daily Ecosystem Summarizer (provides actionable summaries of workspace activities to support informed decisions). This pattern reveals a blind spot: while these needs were voiced by users, they were underestimated by internal practitioners, potentially due to organizational constraints or differing views on feasibility. What is interesting, no feature was overestimated, meaning there were no cases where the market predicted that a feature would be implemented, but it actually was not. This conservative bias suggests that market signals can be trusted to filter out overly speculative ideas (e.g., Brain-computer Interface Coding), but may understate the pace of external innovation.
Alignment with user needs
Among the 12 AI features derived from user needs in Phase 1, three were correctly forecast as likely to be implemented and were in fact released: Multi-file Generation, Project-wide Context Awareness, and Sketch2Code. These features respond to the needs around AI in Implementation and Context-aware Technology. Practitioners identified them as feasible and relevant, and their accurate prediction indicates strong consensus and implementation readiness.
In contrast, features that required deeper system integration or organizational data, such as Proactive Documentation, Automation of Repetitive Tasks, Debugging, Optimization, and AI Real-Time Refactoring, received low probability scores (all below 33%) and were not released. This suggests that despite user interest, practitioners recognized the technical or infrastructural barriers involved. At the same time, the low probabilities assigned to these features do not imply irrelevance. On the contrary, their presence in the prioritized list from Phase 1 and the fact that practitioners recognized them, albeit cautiously, indicates that these needs remain unaddressed by current tools. This creates a clear opportunity for future development.
Our results suggest that prediction markets can help professional communities identify feasible features and areas for innovation, as well as guide product planning and standardization efforts around AI-assisted development.
Answer to RQ3. Practitioners correctly predicted the outcome for 20 of the 24 features, showing that collective judgment aligns with both user needs and actual realization. Features related to implementation and context awareness were accurately anticipated and released. In contrast, features involving proactivity and code maintenance were consistently underestimated, indicating that these needs remain unmet.
V DISCUSSION
This study bridges qualitative insights from developers with internal feasibility assessments, offering a dual-perspective view on in-IDE AI. By integrating user voices and practitioner judgment, we extend prior literature on human–AI experience in software engineering.
While past studies have mapped general expectations for AI tools for coding, our work focuses on in-IDE use, capturing and systematizing nuanced needs shaped by the real-time and context-rich nature of coding environments. Our findings support existing calls for context-aware [14], non-intrusive [32, 17], transparent [33], ethical [13, 12], and customizable [11] AI systems. However, our study also reveals new priorities: integration into broader business workflows, support for skill building, and proactive behavior.
The inclusion of Churners and Non-Users helps surface barriers often missed in adoption-centered research [7]. In particular, there is a need for trust-building mechanisms and accessible onboarding.
Importantly, the prediction market reveals a pattern of conservatism: features that are straightforward to implement are correctly anticipated and realized. In contrast, proactive features and those tied to code maintenance are underestimated despite user interest. This points to a gap between user demand and what teams believe is feasible in the near term.
Together, our two-phase approach illustrates not only what developers want and expect from AI in IDEs, but also which of these expectations are likely to shape product development. Based on our research findings, we propose several key recommendations to guide the future development of in-IDE AI assistants to improve usability, enhance trust, and better integrate AI into developers’ workflows. These recommendations could be generalized to most IDEs and editors since they represent the requests and needs of the users rather than specific features of a specific system.
Technology Improvements
Create Proactive AI. All three participant groups expressed a need for Proactive AI, which was also rated as relevant in the prediction market. AI assistants should take a more proactive role in development, moving beyond reactive responses to explicit prompts and recommendations. Key actions include:
  • Autonomous suggestions. Continuously analyze the code and suggest improvements, such as identifying potential bugs, performance issues, or optimization opportunities.
  • Automated actions. Instead of suggesting changes, the assistant should be able to perform them proactively, e.g., automatically creating new files with boilerplate code or project structure when a user starts a new project.
Build User Trust. Trust is critical for AI adoption in development environments. Developers should prioritize features that build transparency and reliability:
  • On-premise AI. Offer on-premise AI to enhance security and privacy for organizations with sensitive codebases.
  • Transparent outputs. Providing explanations and sources (e.g., Stack Overflow links) for the assistant’s suggestions helps users understand and trust the output.
  • Up-to-date data. Ensure that the AI assistant leverages current and relevant libraries, frameworks, and coding standards to provide reliable recommendations.
Ensure Accessibility. AI technologies should be usable by developers from various regions and backgrounds:
  • Flexible pricing and access. Enable users to choose the features they want to pay for and provide options to access AI from various platforms.
  • Language support. Incorporate multi-language support and localized interfaces to ensure that AI tools are usable in different countries.
  • Inclusive access. Consider regional differences in infrastructure by offering lightweight or offline versions for areas with limited internet connectivity.
Ensure Privacy. Privacy and security were unresolved concerns for many participants. To support adoption under organizational and national constraints, AI systems should:
  • Support on-premise and local AI. Allow use without sending code to external servers, enabling companies to keep data within national or internal infrastructure.
  • Enable auditability. Ensure that AI-generated code is clearly marked to support inspection and maintain trust.
Simplifying Programming Tasks
Look Beyond Code Writing. To support the needs of all users’ groups, AI should not only assist in code writing but also in code-base navigation, evaluation, and in improving code quality:
  • Codebase navigation. Support developers in understanding and exploring the codebase through contextual summaries and links between related components.
  • Project analysis. AI assistants should be able to provide evaluations of the codebase and suggest optimizations for performance, security, or maintainability.
Integrate AI Systems. Participants in all three groups expressed an unmet need for in-IDE AI systems that are tightly connected with internal systems and other engineering tools. Smooth integration with external systems will boost the AI’s utility throughout the SDLC:
  • CI/CD integration. Enable the AI to support continuous integration and continuous deployment pipelines, automating build and deployment tasks.
  • Task tracking integration. Connect with task management tools, allowing the AI to monitor issue statuses, update progress, and link IDE activities with external sources.
  • Commit management. Automate and manage commits, generating commit messages and simplifying the version control process within IDE.
  • Code review support. Integrate with code review platforms to assist in reviewing code, and suggesting improvements without leaving the IDE.
Technology Interaction
Streamline In-IDE Interaction. Participants emphasized the importance of minimizing context switching and ensuring smooth interaction between AI and existing IDE features. To support this:
  • Enable embedded interfaces. Keep AI interaction within the IDE through integrated chat and in-line code explanations to reduce disruption.
  • Support multimodal input. Allow voice commands or device-based control (e.g., wearables) to increase accessibility and flexibility.
Technology Customization
Personalize AI. Developers, especially Churners, seek personalized AI experiences that adapt to their unique needs and workflows. This can be achieved by:
  • Custom setups. Allow users to configure the AI assistant, e.g., selecting models, defining interaction methods, or choosing the technology stack.
  • Learning from interactions. Let the assistant learn from user feedback (e.g., declining specific completions) and adjust its behavior over time.
Simplifying Skill Building
Reduce the Learning Curve. Onboarding and education are key for users to fully utilize AI assistants. All three groups of participants expressed challenges and needs in this regard:
  • User training. Offer tutorials and examples to help users interact effectively with the AI.
  • Prompting assistance. Implement prompt suggestions and provide users with feedback on their prompts to improve their AI interaction skills.
VI THREATS TO VALIDITY
Internal validity. The sample size was determined to reach data saturation; therefore, conclusions are drawn based on the patterns observed within this sample, and further research with a larger or different sample may reveal additional insights or variations of existing topics.
The interview guide was developed and piloted to refine the questions and structure. However, despite these efforts, there may still be inherent biases in the questions or the way they were presented that could influence participants’ responses. The iterative process of developing and refining the interview guide aimed to mitigate this risk.
External validity. The sample was recruited from a pool of developers using JetBrains products. By focusing on JetBrains’ users but allowing for different AI tools, we aimed to balance consistency with a broad range of experiences, ensuring that our findings are applicable beyond any single tool. However, it may not fully capture the diversity of experiences and perspectives of developers using different IDEs, potentially limiting the generalizability of the findings.
Construct validity. Thematic analysis involved iterative coding and discussion sessions among the authors to refine categories and resolve disagreements. Multiple analysts reviewed and coded the data, which helps to ensure consistency and reliability, but some variability in interpretation is possible.
In the second phase, conceptual features derived from interviews were used to formulate prediction market questions. However, these features did not always align well with concrete implementation outcomes. Some definitions remained necessarily broad, which introduced ambiguity into the resolution process. While a structured resolution guide and expert panel were used to mitigate this issue, some interpretive variability may remain, affecting the alignment between predicted outcomes and actual market developments.
VII CONCLUSION
This study examined what developers want and expect from AI in IDEs, focusing on unmet needs and anticipated features. Through interviews with 35 Adopters, Churners, and Non-Users, we surfaced a structured landscape of user expectations across five areas: Technology Improvements, Programming Tasks, Interaction, Customization, and Skill Building. These highlight the demand for more reliable, integrated, and personalized AI experiences within development environments. To assess implementation feasibility, we conducted an internal prediction market with 102 practitioners betting on 12 AI features covering unmet needs. The results show strong alignment between user needs and practitioner judgment for core implementation and context-aware features. However, proactive behaviors and long-term code maintenance remain both underdelivered and underestimated, pointing to a systemic blind spot. Our findings bridge user expectations and product planning by showing where practitioner foresight aligns, or fails to align, with real-world developer needs. Future work should explore how to operationalize unmet needs that are recognized but deprioritized, and how to refine internal evaluation strategies for AI features beyond automation and into workflow-aware design.